# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ 5.2. Structure constants for the equivariant cohomology of $\mbox{Gr}_{\bf b}(k,n)$
Next, we prove the positivity condition of the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$. In [9], the author shows that $\widetilde{{C}}_{i~{}j}^{\ell}$ is a polynomial of $y_{q+1}-y_{q}$ with non negative coefficients for $q=1,2,\dots,n-1$.
@@
The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ can be written as a polynomial of $y_{q}-y_{q+1}-\frac{b_{u}-b_{v}}{b_{0}}Y_{\lambda^{0}}$ with non negative coefficients for $q=1,2,\dots,{n-1}$, where $\lambda^{u}=(q+1,q)\lambda^{v}$ for some $u<v$.
@@
Recall that ${\bf{b}}wt(p_{s})=Y_{\lambda^{e}}-Y_{\lambda^{f}}-\frac{b_{e}-b_{f}}{b_{0}}Y_{\lambda^{0}}$. Let $Y_{\lambda^{e}}-Y_{\lambda^{f}}=y_{r}-y_{s}$ for some $1\leq r<s\leq n$. Then ${\bf{b}}wt(p_{s})$ can be written as the linear sum of the polynomials $(y_{q}-y_{q+1}-\frac{b_{u}-b_{v}}{b_{0}}Y_{\lambda^{0}})$, for $q=r,r+1,\dots,s$. Also $b_{u}-b_{v}>0$ as ${\bf{b}}$ is a divisive Plücker weight vector. Thus, the term $a^{(P)}_{s}$ in the statement of the Theorem 5.14 satisfies the claim. Now we need to prove the same for ${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}$. Note that if $\lambda^{\ell^{\prime}}=(c+1,c)\lambda^{\ell}$ for some $(c+1,c)\in\text{rev}(\lambda^{\ell})$ then $\lambda^{\ell^{\prime}}<\lambda^{\ell}$ and
$(Y_{\lambda^{0}}-\frac{b_{0}}{b_{\ell}}Y_{\lambda^{\ell}})=\frac{b_{\ell^{\prime}}}{b_{\ell}}(Y_{\lambda^{0}}-\frac{b_{0}}{b_{\ell^{\prime}}}Y_{\lambda^{\ell^{\prime}}})+\frac{b_{0}}{b_{\ell}}(y_{c}-y_{c+1}-\frac{b_{\ell^{\prime}}-b_{\ell}}{b_{0}}Y_{\lambda^{0}}).$
Also $Y_{\lambda^{0}}-\frac{b_{0}}{b_{1}}Y_{\lambda^{1}}=\frac{b_{0}}{b_{1}}(Y_{\lambda^{0}}-Y_{\lambda^{1}}-\frac{b_{1}-b_{0}}{b_{0}}Y_{\lambda^{0}})$. This completes the proof. ∎
6. Structure constants for the integral cohomology ring of divisive weighted Grassmann orbifolds
In this section, we compute the integral cohomology rings of divisive weighted Grassmann orbifolds. First, we describe the integral cohomology groups of divisive weighted Grassmann orbifolds using Theorem 4.12. Then, we consider a basis for the cohomology groups and compute the cup product concerning that basis.
@@ 6. Structure constants for the integral cohomology ring of divisive weighted Grassmann orbifolds
For every Schubert symbol $\lambda^{i}$, we can define the weighted Schubert class
${\bf{b}}{\mathbb{S}}_{\lambda^{i}}:=(\xi)^{*}({\bf{b}}{\widetilde{\mathbb{S}}}_{\lambda^{i}}),$
where $\xi$ is the inclusion map
$\xi\colon\mbox{Gr}_{\bf{b}}(k,n)\to ET^{n}\times_{T^{n}}\mbox{Gr}_{\bf{b}}(k,n).$
Geometrically, ${\bf b}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the equivariant cohomology class in the integral equivariant cohomology ring of the divisive weighted Grassmann orbifold corresponding to the closure of the cell $\pi_{\bf b}(\sigma_{r}\widetilde{E}(\sigma_{r}\lambda^{i}))$. Thus $(\xi)^{*}({\bf b}\widetilde{\mathbb{S}}_{\lambda^{i}})$ represents the Poincare dual of the cell $\pi_{\bf b}(\sigma_{r}\widetilde{E}(\sigma_{r}\lambda^{i}))$ in ordinary cohomology. Note that dimension of $\pi_{\bf b}(\sigma_{r}\widetilde{E}(\sigma_{r}\lambda^{i}))$ is $2(k(n-k)-d(\lambda^{i}))$. Therefore, $\{{\bf{b}}{\mathbb{S}}_{\lambda^{i}}:2d(\lambda^{i})=j\}$ forms a basis for the ordinary cohomology group $H^{j}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$.
Let $\lambda^{i}$, $\lambda^{j}$ and $\lambda^{\ell}$ be three Schubert symbols such that $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. Then we have the following product formulae in the ordinary cohomology ring of divisive weighted Grassmann orbifold.
(6.1) ${\bf{b}}{\mathbb{S}}_{\lambda^{i}}{\bf{b}}{\mathbb{S}}_{\lambda^{j}}=\sum_{\lambda^{\ell}~{}:~{}d(\lambda^{\ell})=d(\lambda^{i})+d(\lambda^{j})}{\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}{\bf{b}}{\mathbb{S}}_{\lambda^{\ell}}.$
We make use of Theorem 5.14 to deduce a formulae for the structure constant ${\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}$. Note that the equivariant structure constant ${\bf{b}}{\widetilde{\mathscr{C}}}_{i~{}j}^{\ell}$ coincides with ${\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}$ if $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. We recall for any two Schubert symbol $\lambda^{i}$ and $\lambda^{j}$ we denote $\lambda^{j}\to\lambda^{i}$ if $d(\lambda^{j})=d(\lambda^{i})+1$ and $\lambda^{i}\preceq\lambda^{j}$. We have the following formulae.
@@
Proof.
Proof follows from the proof of Proposition 5.13. ∎
Using Lemma 5.7 it follows $\widetilde{\mathscr{C}}_{i~{}j}^{q}$ is a polynomial of $Y_{\lambda^{e}}-Y_{\lambda^{f}}$ for some $\lambda^{e}\prec\lambda^{f}$. Define
$\widetilde{\mathscr{C}}_{i~{}j}^{q}({\bf{b}}):=\sum_{P~{}:~{}\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\prod_{s=1}^{|P|}{\bf{b}}(p_{s}),$
where ${\bf{b}}(p_{s})$ is defined in (5.5) for every equivariant puzzle pieces $p_{s}$ of $P$ for $s=1,2,\dots,|P|$. For every Schubert symbol $\lambda^{q}$ such that $\lambda^{\ell}\succ\lambda^{q}\succeq\lambda^{i},\lambda^{j}$ and for every $\lambda^{\ell}=\lambda^{\ell_{0}}\rightarrow\lambda^{\ell_{1}}\rightarrow\dots\rightarrow\lambda^{\ell_{d}}=\lambda^{q}$ we can calculate
@@
We recall the formulae for equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ from Theorem 5.14. The maximum value of $d(\lambda^{\ell})$ which satisfy (5.7) is $d(\lambda^{q})+s$. Thus the maximum value of $d(\lambda^{\ell})$ which satisfy (5.2) is $d(\lambda^{q})+|P|$. Now $|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})$. Thus $d(\lambda^{\ell})\leq d(\lambda^{i})+d(\lambda^{j})$. The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ becomes the structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ for the ordinary cohomology if $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. Thus we need to consider $s=|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})=d(\lambda^{\ell})-d(\lambda^{q}).$ Let $d:=d(\lambda^{\ell})-d(\lambda^{q})$.
${\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}$ $=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}a^{(P)}_{d}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{d}}^{\ell}$
$=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{(b_{0})^{d}}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}\prod_{s=1}^{d}(\frac{{\bf{b}}(p_{s})}{b_{0}})$
$=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{1}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\prod_{s=1}^{d}{\bf{b}}(p_{s})$
@@
For a divisive $\mbox{Gr}_{\bf b}(k,n)$ the structure constants ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}\in\mathbb{Z}$ and all are non-negative follows from Theorem 5.15 and Theorem 5.16 respectively.
@@
Consider a Plücker weight vector ${\bf{b}}:=(\alpha\beta,\alpha\beta,\alpha\beta,\alpha,\alpha,\alpha)$, where $\alpha$ and $\beta$ are two positive integers. Then ${\bf{b}}$ is a divisive Plücker weight vector and the corresponding weighted Grassmann orbifold $\mbox{Gr}_{\bf{b}}(2,4)$ is a divisive weighted Grassmann orbifold. In this case, we have $6$ Schubert symbols given by
Corresponding to every Schubert symbol we have $6$ basis element $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{5}$. From Proposition 6.2 it follows that
${\bf{b}}{C}_{1~{}i}^{j}=\frac{b_{0}}{b_{i}}\text{ if }d(\lambda^{j})=d(\lambda^{i})+1\text{ and }\lambda^{j}\succ\lambda^{i}.$
@@
$\widetilde{{S}}_{\lambda^{3}}\widetilde{{S}}_{\lambda^{3}}$ $=(y_{4}-y_{2})(y_{4}-y_{3})\widetilde{{S}}_{\lambda^{3}}+(y_{4}-y_{3})\widetilde{{S}}_{\lambda^{1}}+\widetilde{{S}}_{\lambda^{0}}$
$=(Y_{\lambda^{5}}-Y_{\lambda^{3}})(Y_{\lambda^{2}}-Y_{\lambda^{1}})\widetilde{{S}}_{\lambda^{3}}+(Y_{\lambda^{2}}-Y_{\lambda^{1}})\widetilde{{S}}_{\lambda^{1}}+\widetilde{{S}}_{\lambda^{0}}.$
Applying $\sigma_{r}$ both side we get
@@
From [14]
$\widetilde{{S}}_{\lambda^{3}}\widetilde{{S}}_{\lambda^{2}}=(y_{4}-y_{1})\widetilde{{S}}_{\lambda^{1}}=(Y_{\lambda^{5}}-Y_{\lambda^{1}})\widetilde{{S}}_{\lambda^{1}}.$
Applying $\sigma_{r}$ both side we get
@@
Acknowledgement. The author would like to express my special thanks of gratitude to Soumen Sarkar for his guidance and support. We have many valuable discussions. The author thanks Hiraku Abe and Tommo Matsumura for helpful discussions. The author would like to thank IIT Madras for the Ph.D. fellowship.

### 09 Mar 2026 v3

@@ 4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
Next, we prove the positivity condition of the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ in the sense of [9].
@@
The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ can be written as a polynomial of $y_{r}-y_{r+1}$ for $r=1,2,\dots,{n-1}$ and $(-Y_{\lambda^{0}})$ with coefficients are non negative integers.
@@
The authors in [16] proved that $K(i,j,q,R)\geq 0$. For every $r\in\{1,2,\dots,n\}$ there exist Schubert symbol $\lambda^{e}$ and $\lambda^{f}$ such that $\lambda^{e}<\lambda^{f}$ and $w_{r}-w_{r+1}=b_{e}-b_{f}$. Thus for a divisive weighted Grassmann orbifold $w_{r}-w_{r+1}\geq 0$. From (5.6) it implies that $\mathcal{L}_{s,R}$ is a polynomial of $y_{r}-y_{r+1}$ for $r=1,2,\dots,{n-1}$ with non negative integer coefficients.
$Y_{\lambda^{i}}=(Y_{\lambda^{0}}-Y_{\lambda^{i}})-Y_{\lambda^{0}},\text{ for }i\geq 0.$
Now $(Y_{\lambda^{0}}-Y_{\lambda^{i}})$ can be written as a polynomial in $y_{r}-y_{r+1}$ for $r\in\{1,2,\dots,n-1\}$. Thus ${\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}$ described in (5.9) can be written as a polynomial of $y_{r}-y_{r+1}$ for $r=1,2,\dots,{n-1}$ and $(-Y_{\lambda^{0}})$ with non negative coefficients. Hence the proof. ∎
6. The integral cohomology ring structures of divisive weighted Grassmann orbifolds
In this section, we compute the integral cohomology rings of divisive weighted Grassmann orbifolds. First, we describe the integral cohomology groups of divisive weighted Grassmann orbifolds using Theorem 4.9. Then, we consider a basis for the cohomology groups and compute the cup product concerning that basis. The proof of the following proposition follows from Theorem 4.9 and Corollary 4.10.
@@ 6. The integral cohomology ring structures of divisive weighted Grassmann orbifolds
Let $\varepsilon:\mathbb{Z}[y_{1},\dots,y_{n}]\to\mathbb{Z}$ be the ring homomorphism defined by $\varepsilon(y_{i})=0$. The ordinary cohomology ring $H^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$ can be recovered from the equivariant cohomology ring by base change along $\varepsilon$.
$H^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})\cong H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})\otimes_{\mathbb{Z}[y_{1},\dots,y_{n}]}\mathbb{Z}.$
Geometrically, the map $\varepsilon$ corresponds to trivializing the torus action. Thus, we obtain the forgetful map
$\xi:H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})\longrightarrow H^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z}),$
which simply sends each $y_{i}$ to zero. For every Schubert symbol $\lambda^{i}$, we can define the weighted Schubert class
${\bf{b}}{\mathbb{S}}_{\lambda^{i}}:=\xi({\bf{b}}{\widetilde{\mathbb{S}}}_{\lambda^{i}}).$
Geometrically, ${\bf b}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the equivariant cohomology class in the integral equivariant cohomology ring of the divisive weighted Grassmann orbifold corresponding to the closure of the cell ${E}(\lambda^{i})$ of codimension $2d(\lambda^{i})$. Therefore, $\{{\bf{b}}{\mathbb{S}}_{\lambda^{i}}:2d(\lambda^{i})=j\}$ forms a basis for the ordinary cohomology group $H^{j}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$. Let $\lambda^{i}$, $\lambda^{j}$ and $\lambda^{\ell}$ be three Schubert symbols such that $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. Then we have the following product formulae in the ordinary cohomology ring of divisive weighted Grassmann orbifold.
(6.1) ${\bf{b}}{\mathbb{S}}_{\lambda^{i}}{\bf{b}}{\mathbb{S}}_{\lambda^{j}}=\sum_{\lambda^{\ell}:d(\lambda^{\ell})=d(\lambda^{i})+d(\lambda^{j})}{\bf{b}}{\mathscr{C}}_{i~j}^{\ell}{\bf{b}}{\mathbb{S}}_{\lambda^{\ell}}.$
By the third condition of Definition 5.2, it follows that equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ is a homogeneous polynomial in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ of degree $d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{\ell})$. Applying the forgetful map $\xi$ kills all terms of positive degree in the $y_{i}$, so only those structure constants with $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$ survives. For the special case $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$ the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ are integers and agree with the structure constant ${\bf{b}}{\mathscr{C}}_{i~j}^{\ell}$ in ordinary cohomology. Therefore, Theorem 5.11 yields an explicit formulae for the structure constant ${\bf{b}}{\mathscr{C}}_{i~j}^{\ell}$. We recall for any two Schubert symbol $\lambda^{i}$ and $\lambda^{j}$ we denote $\lambda^{j}\to\lambda^{i}$ if $d(\lambda^{j})=d(\lambda^{i})+1$ and $\lambda^{i}\preceq\lambda^{j}$. We have the following formulae from the Example 5.12.
@@
Define $b(u_{r}):=b_{e}-b_{f}$ for $r\in\{1,2,\dots,n-1\}$ if $y_{r}-y_{r+1}=Y_{\lambda^{e}}-Y_{\lambda^{f}}$. For every finite sub collection $R$ of $\{1,2,\dots,n-1\}$ we define $b(U_{R}):=\prod_{r\in R}b(u_{r})$. Moreover
(6.2) $\widetilde{\mathscr{C}}_{i~j}^{\ell}(b):=\sum_{|R|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{\ell})}K(i,j,\ell,R)b(U_{R})$
For every Schubert symbol $\lambda^{q}$ such that $\lambda^{\ell}\succ\lambda^{q}\succeq\lambda^{i},\lambda^{j}$ and for every sequence $\lambda^{\ell}=\lambda^{\ell_{0}}\rightarrow\lambda^{\ell_{1}}\rightarrow\dots\rightarrow\lambda^{\ell_{d}}=\lambda^{q}$ we can calculate
@@
The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ becomes the structure constant ${\bf{b}}\mathscr{C}_{i~j}^{\ell}$ for the ordinary cohomology if $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. Thus we consider $s=|R|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})=d(\lambda^{\ell})-d(\lambda^{q})$ in (5). Define $d:=d(\lambda^{\ell})-d(\lambda^{q})$.
${\bf{b}}{\mathscr{C}}_{i~j}^{\ell}$ $=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{R:~|R|=d}K(i,j,q,R)\mathcal{L}_{d,R}{\bf{b}}\widetilde{\mathscr{C}}_{q~g^{d}}^{\ell}$
$=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{R:|R|=d}K(i,j,q,R)b(U_{R})\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{1}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}$
@@
We apply Remark 5.10 together with Remark 5.7 to derive the second line from the first line. ∎
@@
For a divisive $\mbox{Gr}_{\bf b}(k,n)$ the structure constants ${\bf{b}}\mathscr{C}_{i~j}^{\ell}\in\mathbb{Z}_{\geq 0}$ follows from Theorem 5.13 together with Theorem 5.14.
@@
Let $\alpha$ and $\beta$ be two positive integers. Consider a Plücker weight vector ${\bf{b}}:=(\alpha\beta,\alpha\beta,\alpha\beta,\alpha,\alpha,\alpha)$ for $k=2,n=4$. Then $\bf{b}$} is a divisive Plücker weight vector and the corresponding weighted Grassmann orbifold $\mbox{Gr}_{\bf{b}}(2,4)$ is a divisive weighted Grassmann orbifold. Corresponding to every Schubert symbol we have $6$ basis element $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{5}$. Using Proposition 6.2, it follows that
${\bf{b}}{\mathscr{C}}_{1~i}^{j}=\frac{b_{0}}{b_{i}}\text{ if }d(\lambda^{j})=d(\lambda^{i})+1\text{ and }\lambda^{j}\succ\lambda^{i}.$
@@
Moreover, using [16]
@@
7. appendix
In this appendix, we discuss some properties on integral cohomology of $\mbox{Gr}_{\bf b}(2,4)$. Recall the subcomplexes $X^{i}$ of $\mbox{Gr}_{\bf{b}}(2,4)$ from (2.14) for $i=0,\dots,5$. Note that
$X^{3}=\mathbb{W}P(b_{5},b_{4},b_{3})\sqcup_{\mathbb{W}P(b_{5},b_{4})}\mathbb{W}P(b_{5},b_{4},b_{2}).\text{ Define }$
$\eta=\frac{\operatorname{lcm}\{b_{5},b_{4},b_{3}\}}{\operatorname{lcm}\{b_{5},b_{4}\}}\text{ and }\eta^{\prime}=\frac{\operatorname{lcm}\{b_{5},b_{4},b_{2}\}}{\operatorname{lcm}\{b_{5},b_{4}\}}.$
Lemma 7.1.
Proof.
Consider the cofibration
$L^{\prime}(b_{3};(b_{5},b_{4}))\to\mathbb{W}P(b_{5},b_{4},b_{2})\to X^{3}.$
This induces an exact sequence in homology by the following
(7.1) $\to\widetilde{H}_{i}(L^{\prime}(b_{3};(b_{5},b_{4})))\to H_{i}(\mathbb{W}P(b_{5},b_{4},b_{2}))\to H_{i}(X^{3})\to\widetilde{H}_{i-1}(L^{\prime}(b_{3};(b_{5},b_{4})))\to$
For $i=1$ in (7.1) we get
$0\to H_{1}(X^{3})\to\mathbb{Z}\to\mathbb{Z}\to.$
$H_{1}(X^{3})$ is either $\{0\}$ or a finite abelian group as $X^{3}$ has no cell of dimension $1$. Also $H_{1}(X^{3})\to\mathbb{Z}$ is injective. Thus $H_{1}(X^{3})=\{0\}$.
For $i=3$ in (7.1), we get
$\to\mathbb{Z}\to 0\to H_{3}(X^{3})\to 0\to\mathbb{Z}.$
Thus $H_{3}(X^{3})=\{0\}$. For $i=4$ in (7.1) we get
$\to 0\to\mathbb{Z}\to H_{4}(X^{3})\to\mathbb{Z}\to 0.$
Thus $H_{4}(X^{3})=\mathbb{Z}\oplus\mathbb{Z}$. For $i\geq 5$ in (7.1) we get
$\to 0\to 0\to H_{i}(X^{3})\to 0\to.$
Now put $i=2$ in (7.1) we get
$\to 0\to\mathbb{Z}\to H_{2}(X^{3})\to\mathbb{Z}_{\eta}\to 0,$
$L^{\prime}(b_{2};(b_{5},b_{4}))\to\mathbb{W}P(b_{5},b_{4},b_{3})\to X^{3}.$
Considering the homology exact sequence corresponding to this cofibration, we can say $H_{2}(X^{3})$ has maximum $\eta^{\prime}$ torsion. Hence, if $\gcd(\eta,\eta^{\prime})=1$ then $H_{i}(X^{3},\mathbb{Z})$ has no torsion for all $i$. ∎
Corollary 7.2.
For a prime $p$, there always exist a Plücker permutation $\sigma$ on $\{0,1,\dots,5\}$ such that the integral cohomology of the 3rd skeleton $X^{3}$ in $\mbox{Gr}_{\sigma\bf b}(2,4)$ is torsion free.
Theorem 7.3.
Proof.
For a prime $p$, choose a Plücker permutation $\sigma$ such that the minimum power of $p$ attains at $b_{\sigma(0)}$. Note that there are $4$ different Plücker permutation which satisfy the above condition. We can choose $\sigma$ such that $r_{\sigma(1)}\leq\min\{r_{\sigma(1)},r_{\sigma(2)},r_{\sigma(3)},r_{\sigma(4)}\}$, where $p^{r_{i}}$ is the $p$-th content in $b_{i}$. Thus using Lemma 3.4, we can assume that $r_{0}\leq r_{1}\leq\min\{r_{1},r_{2},r_{3},r_{4}\}$. From the cofibration $L^{{}^{\prime}}(b_{0};{\bf{b}}^{(0)})\to X^{4}\hookrightarrow\mbox{Gr}_{\bf{b}}(2,4)$ we have
$\to H_{i}(X^{4})\to H_{i}(\mbox{Gr}_{\bf{b}}(2,4))\to\widetilde{H}_{i-1}(L^{\prime}(b_{0};{\bf{b}}^{(0)}))\to H_{i-1}(X^{4})\to.$
For $i=8$, we have
$\to 0\to H_{8}(\mbox{Gr}_{\bf{b}}(2,4))\to\mathbb{Z}\to 0\to.$
For $2\leq i\leq 7$, $\widetilde{H}_{i-1}(L^{\prime}(b_{0};{\bf{b}}^{(0)}))$ has no $p$-torsion. Thus $H_{i}(\mbox{Gr}_{\bf{b}}(2,4))$ has no $p$-torsion if the same is true for $H_{i}(X^{4})$. Similarly from the cofibration $L^{{}^{\prime}}(b_{1},{\bf{b}}^{(1)})\to X^{3}\hookrightarrow X^{4}$ we have
$\to\widetilde{H}_{i}(L^{\prime}(b_{1};{\bf{b}}^{(1)}))\to H_{i}(X^{3})\to H_{i}(X^{4})\to\widetilde{H}_{i-1}(L^{\prime}(b_{1};{\bf{b}}^{(1)}))\to.$
For $i=6$ we
$0\to H_{6}(X^{4})\to\mathbb{Z}\to 0.$
Thus $H_{6}(X^{4})=\mathbb{Z}$. For $2\leq i\leq 5$, $H_{i}(X^{4})$ has no $p$-torsion if the same is true for $H_{i}(X^{3})$ as $\widetilde{H}_{i-1}(L^{\prime}(b_{1};{\bf{b}}^{(1)}))$ has no $p$-torsion. Using Lemma 7.1, $H_{i}(X^{3})$ has no torsion if $i\neq 2$. Thus $H_{i}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion if $i\neq 2$. Hence, the result follows the universal coefficient theorem for cohomology. ∎
Theorem 7.4.
Let ${\bf{b}}=(b_{0},b_{1},b_{2},b_{3},b_{4},b_{5})\in(\mathbb{Z}_{\geq 1})^{6}$ be a Plücker weight vector for $2<4$. Then $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion if $b_{1}=b_{2}$ and $(b_{0},b_{5})=1$.
Proof.
Since $\bf{b}$} is a Plücker weight vector we have $b_{1}+b_{4}=b_{2}+b_{3}$. If $b_{1}=b_{2}$ then $b_{3}=b_{4}$. Thus ${\bf{b}}=(b_{0},b_{1},b_{1},b_{3},b_{3},b_{5})$. For a prime $p$, we can assume $p$-th content of $b_{1}$ divides $p$-th content of $b_{3}$. Otherwise consider the Plücker permutation $\sigma$ defined by $\sigma(1)=4,\sigma(2)=3$, $\sigma(4)=1,\sigma(3)=2$ and $\sigma(i)=i$ for $i\in\{0,5\}$. Then $\sigma{\bf{b}}=(b_{0},b_{3},b_{3},b_{1},b_{1},b_{5})$. Then by Theorem 4.1, $H^{*}(X^{4};\mathbb{Z})$ has no torsion. Considering the cohomology exact sequence in (4.1), $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has maximum $b_{0}$ torsion. Consider the Plücker permutation $\sigma^{\prime}$ defined by $\sigma^{\prime}(0)=5$, $\sigma^{\prime}(5)=0$ and $\sigma^{\prime}(i)=i$ for $i\in\{1,2,3,4\}$. Then $\sigma^{\prime}{\bf{b}}=(b_{5},b_{1},b_{1},b_{3},b_{3},b_{0})$. By a similar argument $H^{*}(\mbox{Gr}_{\sigma^{\prime}{\bf{b}}}(2,4);\mathbb{Z})$ has maximum $b_{5}$ torsion. Since $(b_{0},b_{5})=1$ then $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion. ∎
@@
Acknowledgement. The author thanks to Soumen Sarkar for his guidance and support. We have many valuable discussions. The author thanks Hiraku Abe and Tommo Matsumura for helpful discussions. The author would like to thank IIT Madras for the Ph.D. fellowship.
8. Declarations
Ethical Approval
This declaration is not applicable.
Consent to participate
This declaration is not applicable.
Consent to publish
This declaration is not applicable.
Competing interests
The author declares that he has no competing interests.
Funding
The author is funded by the Indian Institute of Technology Madras PhD fellowship during this work.
Availability of data and materials
Data sharing does not apply to this article, as no data sets were generated or analyzed during the current study.